IQs are designed to be N(100,225).


a. To join Mensa an IQ of 138 is required. What percentage of the population are


eligible to join?


b. A psychologist claims that any child with an IQ of 150+ is ‘gifted’. How many ‘gifted’ children would you expect to find in a school of 1800 pupils?

1
Expert's answer
2022-05-17T15:30:13-0400

a.


"P(X\ge 138)=1-P(Z<\dfrac{138-100}{\sqrt{225}})"

"\approx1-P(Z<2.533)\approx0.00565"

b.


"P(X\ge 150)=1-P(Z<\dfrac{150-100}{\sqrt{225}})"

"\approx1-P(Z<2.533)\approx0.000429"



"0.000429(1800)=1" ‘gifted’ child.